RC meeting and Run Plan, Monday Oct. 9, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary

  • 20:30 - Beam went down

Timeline of events 10/9/23 (Credit Kate)

  • 00:00 - Beam came back around this time and MCC started doing harp scans untill around 01:30
  • 02:27 - Raster parameters were checked. https://logbooks.jlab.org/entry/4195890
  • 02:44 - Target moved from No Target to Carbon Hole. Beam current ~5uA
  • 02:49 - Beam position alarm went off. Beam position off ideal plot read 2.02mm, but we were on carbon hole at this time.
  • 03:19 - Pumping chamber temperature dropped from 285 degrees to 249 degrees within 2 minutes. No alarm went off (LO Setpoint is 180), and the shift crew didn't notice because they were working on beam checkout. The beam current and beam position didn't noticeably change at this time. We suspect something happened at this time.
  • 03:21 - Raster was checked again, coincidentally at about the same time as the drop in PC temperature. https://logbooks.jlab.org/entry/4195912
  • 04:16 - Carbon hole results are posted. https://logbooks.jlab.org/entry/4195935 https://logbooks.jlab.org/entry/4195936
  • 04:22 - Moved to No Target. Positions looked good with beam. SOFs were properly changed to He3 target settings.
  • 04:38 - Moved to Pickup Coil target position.
  • 04:49 - 3 NMR measurements were taken with results all less than 2% https://logbooks.jlab.org/entry/4195953
  • 04:50 - Moved to Pol He 3 cell target position. Beam position looked good. Tranfer tube temperatures had seen no obvious change during this time.
  • Start of day shift, hall began preparing to go to controlled access. Magnets were ramped down, lasers locked out and hall swept and HV locked out.
  • 12:00 - Escorted access to view inside target enclosure confirmed that the cell was ruptured.

Accelerator status

  • Machine was down 20:30 - ~00:00 for 1L10 dropped out of HV due to breaker issue.
  • There will be a maintenance day during Day shift on Wednesday 10/11
  • If you want to get into the hall, please communicate with the RC to get this organized!

Data collected :

  • DAY 0.740 C
  • SWING: 0.207 C
  • OWL: N/A
  • DAY: N/A
  • TOTAL: 0.947 (Sunday day up till rupture.)

Hall A Subsystems Status


  • Ruptured. See summary above.


  • All HV currently off. Hall is in controlled access.
  • BBCal 7.4 dead. Discussion on changing PMT base at next available time.


  • SBS and BB are off. Hall is in controlled access.

Hall access requests

If anyone wants to make an access, let the RC know!

Shift schedule

Please check when your shifts are!


Please sign up for any open shifts!

Running Plan

  • Discuss clean up steps
  • Potential for moller measurement in the evening
  • Cell installation plan tuesday/wednesday - target group.

When we are stable in a few days

  • Standard production running at 45uA, with NMRs every 3 hours.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)